# ToMCCA-3: A realistic 3-body coalescence model

### v1

@@ 2.1 Generalized 3-body coalescence approach
As discussed before, the wave function for the ${{}^{3}_{\Lambda}{\rm H}}$ nucleus can be written in terms of the center-of-mass motion and an internal wave function as
$\Psi_{NR}\left({\bm{x}}_{1},{\bm{x}}_{2},{\bm{x}}_{3}\right)=\left(2\pi\right)^{-3/2}e^{i{\bm{P}}_{{}^{3}_{\Lambda}\text{H}}\cdot{\bm{R}}}\varphi_{{{}^{3}_{\Lambda}\text{H}}}(\bm{\xi}_{1},\bm{\xi}_{2})\ ,$ (23)
where $\varphi_{{}^{3}_{\Lambda}\text{H}}$ is the internal part of the ${{}^{3}_{\Lambda}{\rm H}}$ wave function, describing the dynamics of the p, n, and $\Lambda$ and $\bm{\xi}_{1}$ and $\bm{\xi}_{2}$ are the hyperspherical coordinates defined above. In Eq. (6) changing the integration variables from the particle positions and momenta to the center-of-mass and Jacobi coordinates , one obtains
$\frac{\mathrm{d}N_{{{}^{3}_{\Lambda}{\rm H}}}}{\mathrm{d}^{3}P}$ $=$ $S_{\mathrm{{{}^{3}_{\Lambda}{\rm H}}}}{|J^{-1}|^{3}\over(2\pi)^{15}}\int\mathrm{d}^{3}R\,\mathrm{d}^{3}\xi_{1}\,\mathrm{d}^{3}\xi_{2}\int\mathrm{d}^{3}R^{\prime}\,\mathrm{d}^{3}\,\xi_{1}^{\prime}\mathrm{d}^{3}\xi_{2}^{\prime}\,e^{i\mathbf{P}_{{}^{3}_{\Lambda}\text{H}}\cdot({\bm{R}}-{\bm{R}}^{\prime})}\varphi_{{}^{3}_{\Lambda}\text{H}}^{*}(\bm{\xi}_{1}^{\prime},\bm{\xi}_{2}^{\prime})\varphi_{{}^{3}_{\Lambda}\text{H}}(\bm{\xi}_{1},\bm{\xi}_{2})$ (24)
×$\times\int\mathrm{d}^{3}P\,\mathrm{d}^{3}k_{1}\,\mathrm{d}^{3}k_{2}\,e^{-i\mathbf{P}\cdot\left({\bm{R}}-{\bm{R}}^{\prime}\right)-i{\bm{k}}_{1}\cdot\left(\bm{\xi}_{1}-\bm{\xi}_{1}^{\prime}\right)-i{\bm{k}}_{2}\cdot\left(\bm{\xi}_{2}-\bm{\xi}_{2}^{\prime}\right)}$
×$\times{\tau^{3\over 2}\over(2\pi\sigma^{2})^{9\over 2}}e^{-{2+\tau\over 8\sigma^{2}}\bigl{(}({\bm{R}}+{\bm{R}}^{\prime})^{2}+2({\bm{R}}+{\bm{R}}^{\prime})\cdot{\bm{t}}\bigr{)}}\;e^{-{1\over 16\sigma^{2}}\bigl{(}\alpha(\bm{\xi}_{1}+\bm{\xi}_{1}^{\prime})^{2}+(\bm{\xi}_{2}+\bm{\xi}_{2}^{\prime})^{2}\bigr{)}}G_{\mathrm{np}\Lambda}(\bm{p}_{1},\mathbf{p}_{2},\bm{p}_{3})\ ,$
where $|J^{-1}|=\sqrt{(2+\kappa)/4\kappa}$. Furthermore, by substituting ${\bm{R}}$ and ${\bm{R}}^{\prime}$ with the variables ${\bm{Z}}={\bm{R}}+{\bm{R}}^{\prime}$ and ${\bm{Z}}^{\prime}={\bm{R}}-{\bm{R}}^{\prime}$, the integration over ${\bm{Z}}^{\prime}$ results in a factor $(2\pi)^{3}\delta({\bm{P}}-{\bm{P}}_{{}^{3}_{\Lambda}\text{H}})$, while the integral over ${\bm{Z}}$ results in
$\int d^{3}Z\;e^{-{2+\tau\over 8\sigma^{2}}[Z^{2}+2{\bm{t}}\cdot{\bm{Z}}]}=\int d^{3}Z\;e^{-{2+\tau\over 8\sigma^{2}}[({\bm{Z}}+{\bm{t}})^{2}-t^{2}]}=\left({8\pi\sigma^{2}\over 2+\tau}\right)^{3\over 2}e^{{2+\tau\over 8\sigma^{2}}t^{2}}\ .$ (25)
@@
$\frac{\mathrm{d}N_{{{}^{3}_{\Lambda}{\rm H}}}}{\mathrm{d}^{3}P}$ $=$ $S_{\mathrm{{{}^{3}_{\Lambda}{\rm H}}}}{|J^{-1}|^{3}\over(2\pi)^{12}}\int\mathrm{d}^{3}\xi_{1}\,\mathrm{d}^{3}\xi_{2}\int\mathrm{d}^{3}\,\xi_{1}^{\prime}\mathrm{d}^{3}\xi_{2}^{\prime}\,\varphi_{{}^{3}_{\Lambda}\text{H}}^{*}(\bm{\xi}_{1}^{\prime},\bm{\xi}_{2}^{\prime})\varphi_{{}^{3}_{\Lambda}\text{H}}(\bm{\xi}_{1},\bm{\xi}_{2})$ (26)
×$\times\int\,\mathrm{d}^{3}k_{1}\,\mathrm{d}^{3}k_{2}\,e^{-i{\bm{k}}_{1}\cdot\left(\bm{\xi}_{1}-\bm{\xi}_{1}^{\prime}\right)-i{\bm{k}}_{2}\cdot\left(\bm{\xi}_{2}-\bm{\xi}_{2}^{\prime}\right)}$
×$\times{\tau^{3\over 2}\over(2\pi\sigma^{2})^{9\over 2}}e^{{2+\kappa\over 8\sigma^{2}}t^{2}}\;e^{-{1\over 16\sigma^{2}}\bigl{(}\alpha(\bm{\xi}_{1}+\bm{\xi}_{1}^{\prime})^{2}+(\bm{\xi}_{2}+\bm{\xi}_{2}^{\prime})^{2}\bigr{)}}G_{\mathrm{np}\Lambda}(\bm{p}_{1},\bm{p}_{2},\mathbf{p}_{3})|_{{\bm{P}}={\bm{P}}_{{}^{3}_{\Lambda}\text{H}}}\ .$
@@
${2+\tau\over 8\sigma^{2}}t^{2}-{\alpha\over 16\sigma^{2}}(\bm{\xi}_{1}+\bm{\xi}_{1}^{\prime})^{2}={\alpha^{\prime}\over 16\sigma^{2}}(\bm{\xi}_{1}+\bm{\xi}_{1}^{\prime})^{2}\ ,\qquad\alpha^{\prime}={\tau(2+\kappa)\over\kappa(2+\tau)}\ .$ (27)
This is useful to define a compact form of the expression, which allows for the evaluation of the probability of cluster formations as given below.
$\frac{\mathrm{d}N_{{{}^{3}_{\Lambda}{\rm H}}}}{\mathrm{d}^{3}P}={S_{\mathrm{{{}^{3}_{\Lambda}{\rm H}}}}\over(2\pi)^{12}}\int\mathrm{d}^{3}k_{1}\,\mathrm{d}^{3}k_{2}\,G_{np\Lambda}(\mathbf{p}_{1},\mathbf{p}_{2},\mathbf{p}_{3})|_{{\bm{P}}={\bm{P}}_{H}}\;{\cal P}({\bm{k}}_{1},{\bm{k}}_{2},\sigma,\sigma^{\prime})\ ,$ (28)
@@
${\cal P}({\bm{k}}_{1},{\bm{k}}_{2},\sigma,\sigma^{\prime})$ $=$ ${(\alpha^{\prime})^{3\over 2}\over 8(2\pi)^{3}\sigma^{6}}\int\mathrm{d}^{3}\xi_{1}\,\mathrm{d}^{3}\xi_{2}\int\mathrm{d}^{3}\,\xi_{1}^{\prime}\mathrm{d}^{3}\xi_{2}^{\prime}\;\varphi_{{}^{3}_{\Lambda}\text{H}}^{*}(\bm{\xi}_{1}^{\prime},\bm{\xi}_{2}^{\prime})\varphi_{{}^{3}_{\Lambda}\text{H}}(\bm{\xi}_{1},\bm{\xi}_{2})$ (29)
×$\times e^{-i{\bm{k}}_{1}\cdot\left(\bm{\xi}_{1}-\bm{\xi}_{1}^{\prime}\right)-i{\bm{k}}_{2}\cdot\left(\bm{\xi}_{2}-\bm{\xi}_{2}^{\prime}\right)}\;e^{-{1\over 16\sigma^{2}}\bigl{(}\alpha^{\prime}(\bm{\xi}_{1}+\bm{\xi}_{1}^{\prime})^{2}+(\bm{\xi}_{2}+\bm{\xi}_{2}^{\prime})^{2}\bigr{)}}\ .$
Using realistic numbers and Eq. 27, we can approximate $\alpha^{\prime}$ such that
[…]
$\tau$ $=(\sigma_{p}/\sigma_{\Lambda})^{2}\simeq 0.608\Rightarrow\alpha^{\prime}\simeq 0.626$ (31)
As a test, ${\cal P}(\bm{0},\bm{0},\sigma\rightarrow 0,\sigma^{\prime}=\sigma\sqrt{\tau})$ can be computed. In such a case, the Gaussians become delta functions, namely
@@
A complete description of the calculation of $\varphi_{{}^{3}_{\Lambda}\text{H}}$ using the PHH method is provided in the Appendix. B. The case of 3He and 3H can be recovered by setting $\kappa=1$ and $\tau=1$, from which follows that $\alpha^{\prime}=1$ in Eq. 29. Details of the 3He and 3H wave functions can also be found in Appendix B.
@@ 2.2 Congleton approach for Hypertriton
A second approach for calculating the ${}^{3}_{\Lambda}\text{H}$ yield has been developed based on the Congleton wave function [41]. This is a relatively simple approach for determining the ${}^{3}_{\Lambda}\text{H}$ wave function, where the $\Lambda$ particle is assumed to orbit an unperturbed deuteron in a $\Lambda$–d potential ($V_{\Lambda-\mathrm{d}}$). This potential is based on a separable N–$\Lambda$ interaction. The governing Hamiltonian for the system, including center-of-mass motion, is given by
@@
The Hamiltonian is split into two parts using a static approximation, in which the effective $\Lambda$–d system is considered. This way, the influence of the $\Lambda$ particle on the individual nucleons is neglected. Consequently, the $V_{\Lambda-\mathrm{d}}$ potential depends only on the variable $\boldsymbol{q}$ (after accounting for spin dependence). Furthermore, the reduced mass of the $\Lambda$–d system is $\mu=3.547\,\mathrm{fm}^{-1}$ while the reduced mass of the two-nucleon system is $m_{\mu}=2.379\,\mathrm{fm}^{-1}$.
The wave function can be expressed as the product of the spin function ($s=\frac{1}{2}$ for Hypertriton) and the nuclear wave function part
@@
where $\psi_{\Lambda}(q)$ and $\psi_{\mathrm{d}}^{(l)}(p)$ represent the $\Lambda$ wave function and the radial part of the deuteron wave function in the momentum space, respectively. The spin component describes the coupling between the spin-1 deuteron and the spin-$\frac{1}{2}$ $\Lambda$ particle, resulting in a total spin of either $\frac{1}{2}$ or $\frac{3}{2}$. The $s$-wave component of the deuteron corresponds to $S=\frac{1}{2}$, while the $d$-wave component corresponds to $S=\frac{3}{2}$. Apart from the spin component, the spatial wave function is a simple product of the deuteron and $\Lambda$ wave functions. Similar to the HH approach, the total wave function for the hypertriton $\Psi_{{{}^{3}_{\Lambda}{\rm H}}}\left({\bm{x}}_{1},{\bm{x}}_{2},{\bm{x}}_{3}\right)$ can be factorized into a plane wave for the center-of-mass (COM) motion and an internal wave function that depends only on two relative coordinates, as shown in Eq. 23. The internal wave function is given as the product of $\Lambda$ and deuteron wave functions.
@@
$\mathcal{P}(\bm{k}_{1},\bm{k}_{2},\sigma)=\frac{\left(2+\kappa^{2}\right)^{3/2}}{(2+\kappa)^{3}}\ \frac{S_{\mathrm{{{}^{3}_{\Lambda}{\rm H}}}}}{(2\pi)^{3}\sigma^{6}}\,\,\int\mathrm{d}^{3}r_{1}\int\mathrm{d}^{3}r_{2}\,\,\mathcal{D}_{\Lambda}({\bm{k}}_{1},\mathbf{r}_{1})\mathcal{D}_{d}(\bm{k}_{2},\mathbf{r}_{2})e^{-\frac{r_{1}^{2}(2+\kappa)^{2}+4r_{2}^{2}\left(2+\kappa^{2}\right)}{4\sigma^{2}(2+\kappa)^{2}}}.$ (37)
The Wigner densities of the $\Lambda$ particle denoted as $\mathcal{D}_{\Lambda}({\bm{k}}_{1},\bm{r}_{1})$ is calculated in momentum space and the deuteron one, $\mathcal{D}_{d}(\bm{k}_{2},\bm{r}_{2})$, is computed in coordinate space. The computation of $\mathcal{D}_{d}(\bm{k}_{2},\bm{r}_{2})$ has already been provided in a previous work [27], using the Argonne $\text{Argonne }v_{18}$ wave function, while the derivation of $\mathcal{D}_{\Lambda}({\bm{k}}_{1},\bm{r}_{1})$ is given in Appendix C. The radial $\Lambda$ wave function in momentum space is then given by [41]
$\psi_{\Lambda}(q)=N(Q_{\Lambda})\frac{e^{-\left(\frac{q}{Q_{\Lambda}}\right)^{2}}}{q^{2}+\alpha_{\Lambda}^{2}},$ (38)
@@
The Cerfe function is defined as $\mathrm{Cerfe}(x)=\exp(x^{2})(1-\mathrm{erf}(x))$. In the original publication [41] the parameter $\alpha_{\Lambda}=\sqrt{\mu B_{\Lambda}}=0.068$ fm-1 was fixed to the binding energy and $Q_{\Lambda}=1.17$ fm-1 was adjusted to the $\Lambda$ separation energy. In Ref. [24], the parameter $Q_{\Lambda}$ of the wave function was adjusted to match the form factor of Hildenbrand&Hammer [42] obtained from pionless effective field theory. An excellent agreement is reached when $Q_{\Lambda}=2.5$ fm-1 instead of $Q_{\Lambda}=1.17$ fm-1. With these parameters $N(Q_{\Lambda}=2.5\leavevmode\nobreak\ \mathrm{fm}^{-1})^{2}=0.094$ and $\int|\psi_{\Lambda}(q)|^{2}\mathrm{d}^{3}q=4\pi$. The final Wigner function $\mathcal{D}_{\Lambda}(\mathbf{q},\mathbf{r})$ is given as
$\mathcal{D}_{\Lambda}(\mathbf{q},\mathbf{r})=N(Q_{\Lambda})^{2}\int_{0}^{\infty}\mathrm{d}k\frac{4i\pi\left(-1+e^{2ikr}\right)}{qr\left(k^{2}+4\left(\alpha_{\Lambda}^{2}+q^{2}\right)\right)}\log\left(\frac{4\alpha_{\Lambda}^{2}+(k-2q)^{2}}{4\alpha_{\Lambda}^{2}+(k+2q)^{2}}\right)e^{-\frac{k^{2}+2ikQ_{\Lambda}^{2}r+4q^{2}}{2Q_{\Lambda}^{2}}},$ (40)

### v2

@@ 2.1 Generalized 3-body coalescence approach
As discussed before, the wave function for the $A=3$ nucleus can be written in terms of the center-of-mass motion and an internal wave function as
$\Psi_{NR}\left({\bm{x}}_{1},{\bm{x}}_{2},{\bm{x}}_{3}\right)=\left(2\pi\right)^{-3/2}e^{i{\bm{P}}_{3}\cdot{\bm{R}}}\varphi_{{3}}(\bm{\xi}_{1},\bm{\xi}_{2})\ ,$ (21)
where $\varphi_{3}$ is the internal part of the $A=3$ wave function, describing the dynamics of the p, n, and the third baryon and $\bm{\xi}_{1}$ and $\bm{\xi}_{2}$ are the Jacobi coordinates defined above. In Eq. (4) changing the integration variables from the particle positions and momenta to the center-of-mass and Jacobi coordinates, one obtains
$\frac{\mathrm{d}N_{3}}{\mathrm{d}^{3}P}$ $=$ $S_{3}{|J^{-1}|^{3}\over(2\pi)^{15}}\int\mathrm{d}^{3}\bm{R}\,\mathrm{d}^{3}\bm{\xi}_{1}\,\mathrm{d}^{3}\bm{\xi}_{2}\int\mathrm{d}^{3}\bm{R}^{\prime}\,\mathrm{d}^{3}\,\bm{\xi}_{1}^{\prime}\mathrm{d}^{3}\bm{\xi}_{2}^{\prime}\,\exp\left({i\bm{P}_{3}\cdot({\bm{R}}-{\bm{R}}^{\prime})}\varphi_{3}^{*}(\bm{\xi}_{1}^{\prime},\bm{\xi}_{2}^{\prime})\varphi_{3}(\bm{\xi}_{1},\bm{\xi}_{2})\right)$ (22)
$\times\int\mathrm{d}^{3}\bm{P}\,\mathrm{d}^{3}\bm{k}_{1}\,\mathrm{d}^{3}\bm{k}_{2}\,\exp\left({-i\bm{P}\cdot\left({\bm{R}}-{\bm{R}}^{\prime}\right)-i{\bm{k}}_{1}\cdot\left(\bm{\xi}_{1}-\bm{\xi}_{1}^{\prime}\right)-i{\bm{k}}_{2}\cdot\left(\bm{\xi}_{2}-\bm{\xi}_{2}^{\prime}\right)}\right)$
$\times{\tau^{3\over 2}\over(2\pi\sigma^{2})^{9\over 2}}e^{-{2+\tau\over 8\sigma^{2}}\bigl(({\bm{R}}+{\bm{R}}^{\prime})^{2}+2({\bm{R}}+{\bm{R}}^{\prime})\cdot{\bm{t}}\bigr)}\;\exp\left({-{1\over 16\sigma^{2}}\bigl(\alpha(\bm{\xi}_{1}+\bm{\xi}_{1}^{\prime})^{2}+(\bm{\xi}_{2}+\bm{\xi}_{2}^{\prime})^{2}\bigr)}\right)$
$\times G_{3}(\bm{p}_{1},\bm{p}_{2},\bm{p}_{3})\ ,$
where $|J^{-1}|=\sqrt{(2+\kappa)/4\kappa}$. Furthermore, by substituting $\bm{R}$} and $\bm{R}$$^{\prime}$ with the variables ${\bm{Z}}={\bm{R}}+{\bm{R}}^{\prime}$ and ${\bm{Z}}^{\prime}={\bm{R}}-{\bm{R}}^{\prime}$, the integration over $\bm{Z}$$^{\prime}$ results in a factor $(2\pi)^{3}\delta({\bm{P}}-{\bm{P}}_{3})$, while the integral over $\bm{Z}$} results in
$\int\mathrm{d}^{3}\bm{Z}\;\exp\left(-{2+\tau\over 8\sigma^{2}}[Z^{2}+2{\bm{t}}\cdot{\bm{Z}}]\right)=\int\mathrm{d}^{3}\bm{Z}\;\exp\left(-{2+\tau\over 8\sigma^{2}}[({\bm{Z}}+{\bm{t}})^{2}-t^{2}]\right)=\left({8\pi\sigma^{2}\over 2+\tau}\right)^{3\over 2}\exp\left({2+\tau\over 8\sigma^{2}}t^{2}\right)\ .$ (23)
@@
$\frac{\mathrm{d}N_{3}}{\mathrm{d}^{3}P}$ $=$ ${S_{3}\over 8}{|J^{-1}|^{3}\over(2\pi)^{12}}\int\mathrm{d}^{3}\bm{\xi}_{1}\,\mathrm{d}^{3}\bm{\xi}_{2}\int\mathrm{d}^{3}\,\bm{\xi}_{1}^{\prime}\mathrm{d}^{3}\bm{\xi}_{2}^{\prime}\,\varphi_{3}^{*}(\bm{\xi}_{1}^{\prime},\bm{\xi}_{2}^{\prime})\varphi_{3}(\bm{\xi}_{1},\bm{\xi}_{2})$ (24)
$\times\int\,\mathrm{d}^{3}\bm{k}_{1}\,\mathrm{d}^{3}\bm{k}_{2}\,e^{-i{\bm{k}}_{1}\cdot\left(\bm{\xi}_{1}-\bm{\xi}_{1}^{\prime}\right)-i{\bm{k}}_{2}\cdot\left(\bm{\xi}_{2}-\bm{\xi}_{2}^{\prime}\right)}$
$\times{\tau^{3\over 2}\over(2\pi\sigma^{2})^{9\over 2}}\,\left({8\pi\sigma^{2}\over 2+\tau}\right)^{3\over 2}\exp\left({2+\tau\over 8\sigma^{2}}t^{2}\right)\;\exp\left(-{1\over 16\sigma^{2}}\bigl(\alpha(\bm{\xi}_{1}+\bm{\xi}_{1}^{\prime})^{2}+(\bm{\xi}_{2}+\bm{\xi}_{2}^{\prime})^{2}\bigr)\right)$
$\times G_{3}(\bm{p}_{1},\bm{p}_{2},\bm{p}_{3})|_{{\bm{P}}={\bm{P}}_{3}}\ .$
@@
${2+\tau\over 8\sigma^{2}}t^{2}-{\alpha\over 16\sigma^{2}}(\bm{\xi}_{1}+\bm{\xi}_{1}^{\prime})^{2}=-{\alpha^{\prime}\over 16\sigma^{2}}(\bm{\xi}_{1}+\bm{\xi}_{1}^{\prime})^{2}\ ,\qquad\alpha^{\prime}={\tau(2+\kappa)\over\kappa(2+\tau)}\ .$ (25)
This is useful to define a compact form of the expression, which allows for the evaluation of the probability of cluster formation as given below. In the ase of the hypertriton, such probability reads
$\frac{\mathrm{d}N_{{{}^{3}_{\Lambda}{\rm H}}}}{\mathrm{d}^{3}P}={S_{3}\over(2\pi)^{12}}\int\mathrm{d}^{3}\bm{k}_{1}\,\mathrm{d}^{3}\bm{k}_{2}\,G_{3}(\bm{p}_{1},\bm{p}_{2},\bm{p}_{3})|_{{\bm{P}}={\bm{P}}_{3}}\;{\cal P}({\bm{k}}_{1},{\bm{k}}_{2},\sigma,\sigma^{\prime})\ ,$ (26)
@@
${\cal P}({\bm{k}}_{1},{\bm{k}}_{2},\sigma,\sigma^{\prime})$ $=$ ${(\alpha^{\prime})^{3\over 2}\over 8(2\pi)^{3}\sigma^{6}}\int\mathrm{d}^{3}\bm{\xi}_{1}\,\mathrm{d}^{3}\bm{\xi}_{2}\int\mathrm{d}^{3}\,\bm{\xi}_{1}^{\prime}\mathrm{d}^{3}\bm{\xi}_{2}^{\prime}\;\varphi_{{{}^{3}_{\Lambda}{\rm H}}}^{*}(\bm{\xi}_{1}^{\prime},\bm{\xi}_{2}^{\prime})\varphi_{{{}^{3}_{\Lambda}{\rm H}}}(\bm{\xi}_{1},\bm{\xi}_{2})$ (27)
$\times e^{-i{\bm{k}}_{1}\cdot\left(\bm{\xi}_{1}-\bm{\xi}_{1}^{\prime}\right)-i{\bm{k}}_{2}\cdot\left(\bm{\xi}_{2}-\bm{\xi}_{2}^{\prime}\right)}\;e^{-{1\over 16\sigma^{2}}\bigl(\alpha^{\prime}(\bm{\xi}_{1}+\bm{\xi}_{1}^{\prime})^{2}+(\bm{\xi}_{2}+\bm{\xi}_{2}^{\prime})^{2}\bigr)}\ .$
Eq. (27) is finally used inside of ToMCCA. It is evaluated on a 4D $(|\textbf{k}_{1}|,|\textbf{k}_{2}|,\cos{\theta_{k_{12}}},\sigma)$ grid with a predetermined $\alpha^{\prime}$. This evaluation is done once in advance, and the grid is saved in a 4D histogram. As a test, ${\cal P}(\bm{0},\bm{0},\sigma\rightarrow 0,\sigma^{\prime}=\sigma\sqrt{\tau})$ can be computed. In such a case, the Gaussians become delta functions, namely
@@
The case of 3He and 3H can be recovered by setting $\kappa=1$ and $\tau=1$, from which follows that $\alpha^{\prime}=1$ in Eq. (27). Details of the 3He and 3H wave functions can be found in Appendix F.
@@ 2.2 Congleton approach for Hypertriton
[…]
The governing Hamiltonian for the system, including the center-of-mass motion, is given by
@@
The Hamiltonian is split into two parts using a static approximation, in which the effective $\Lambda$–d system is considered. This way, the influence of the $\Lambda$ particle on the individual nucleons is neglected. Consequently, the $V_{\Lambda-\mathrm{d}}$ potential depends only on the $\Lambda$ momentum in the center of mass frame $\boldsymbol{q}$ (after accounting for spin dependence). Furthermore, the reduced mass of the $\Lambda$–d system is $\mu=3.547\,\mathrm{fm}^{-1}$ while the reduced mass of the two-nucleon system is $m_{\mu}=2.379\,\mathrm{fm}^{-1}$. The wave function can be expressed as the product of the spin function ($s=\frac{1}{2}$ for Hypertriton) and the nuclear wave function part
@@
where $\psi_{\Lambda}(q)$ and $\psi_{\mathrm{d}}^{(l)}(p)$ represent the $\Lambda$ wave function and the radial part of the deuteron wave function in the momentum space, respectively. The spin component describes the coupling between the spin-1 deuteron and the spin-$\frac{1}{2}$ $\Lambda$ particle, resulting in a total spin of either $\frac{1}{2}$ or $\frac{3}{2}$. The $s$-wave component of the deuteron corresponds to $S=\frac{1}{2}$, while the $d$-wave component corresponds to $S=\frac{3}{2}$. Apart from the spin component, the spatial wave function is a simple product of the deuteron and $\Lambda$ wave functions. Similar to the HH approach, the total wave function for the hypertriton $\Psi_{{{}^{3}_{\Lambda}{\rm H}}}\left({\bm{x}}_{1},{\bm{x}}_{2},{\bm{x}}_{3}\right)$ can be factorized into a plane wave for the center-of-mass motion and an internal wave function that depends only on two relative coordinates, as shown in Eq. 21. The internal wave function is given as the product of $\Lambda$ and deuteron wave functions.
@@
$\mathcal{P}(\bm{k}_{1},\bm{k}_{2},\sigma)=\frac{\left(2+\kappa^{2}\right)^{3/2}}{(2+\kappa)^{3}}\ \frac{S_{\mathrm{{{}^{3}_{\Lambda}{\rm H}}}}}{(2\pi)^{3}\sigma^{6}}\,\,\int\mathrm{d}^{3}\bm{r}_{1}\int\mathrm{d}^{3}\bm{r}_{2}\,\,\mathcal{D}_{\Lambda}({\bm{k}}_{1},\bm{r}_{1})\mathcal{D}_{d}(\bm{k}_{2},\bm{r}_{2})\exp\left(-\frac{r_{1}^{2}(2+\kappa)^{2}+4r_{2}^{2}\left(2+\kappa^{2}\right)}{4\sigma^{2}(2+\kappa)^{2}}\right).$ (33)
The Wigner densities of the $\Lambda$ particle denoted as $\mathcal{D}_{\Lambda}({\bm{k}}_{1},\bm{r}_{1})$ is calculated in momentum space and the deuteron one, $\mathcal{D}_{d}(\bm{k}_{2},\bm{r}_{2})$, is computed in coordinate space. The computation of $\mathcal{D}_{d}(\bm{k}_{2},\bm{r}_{2})$ has already been provided in a previous work [41], using the $\text{Argonne }v_{18}$ wave function, while the derivation of $\mathcal{D}_{\Lambda}({\bm{k}}_{1},\bm{r}_{1})$ is given in Appendix G. The radial $\Lambda$ wave function in momentum space is then given by [34]
$\psi_{\Lambda}(q)=N(Q_{\Lambda})\frac{\exp\left(-\left(\frac{q}{Q_{\Lambda}}\right)^{2}\right)}{q^{2}+\alpha_{\Lambda}^{2}},$ (34)
@@
The Cerfe function is defined as $\mathrm{Cerfe}(x)=\exp(x^{2})(1-\mathrm{erf}(x))$. In the original publication [34] the parameter $\alpha_{\Lambda}=\sqrt{\mu B_{\Lambda}}=0.068$ fm-1 was fixed to the binding energy and $Q_{\Lambda}=1.17$ fm-1 was adjusted to the $\Lambda$ separation energy. In Ref. [23], the parameter $Q_{\Lambda}$ of the wave function was adjusted to match the form factor of Ref. [35] obtained from $\cancel{\pi}$EFT. An excellent agreement is reached when $Q_{\Lambda}=2.5$ fm-1 instead of $Q_{\Lambda}=1.17$ fm-1. With these parameters $N(Q_{\Lambda}=2.5\penalty 10000\ \mathrm{fm}^{-1})^{2}=0.094369$ fm-1 and $\int|\psi_{\Lambda}(q)|^{2}\mathrm{d}^{3}q=4\pi$. The final Wigner function $\mathcal{D}_{\Lambda}(\bm{q},\bm{r})$ is given as
$\mathcal{D}_{\Lambda}(\bm{q},\bm{r})=N(Q_{\Lambda})^{2}\int_{0}^{\infty}\mathrm{d}k\frac{4i\pi\left(-1+e^{2ikr}\right)}{qr\left(k^{2}+4\left(\alpha_{\Lambda}^{2}+q^{2}\right)\right)}\log\left(\frac{4\alpha_{\Lambda}^{2}+(k-2q)^{2}}{4\alpha_{\Lambda}^{2}+(k+2q)^{2}}\right)\exp\left(-\frac{k^{2}+2ikQ_{\Lambda}^{2}r+4q^{2}}{2Q_{\Lambda}^{2}}\right),$ (36)